OpenStudy (anonymous):

write a polynomial function f of least degree that has rational coefficients, a leading coefficient of 1 , and the zeros 2, -3, and -3i

OpenStudy (anonymous):

Try this: leading coff. a=1 Poly. 1(x-2)(x+3)(x+3i)=0 Multiply them together: (x^2+x-6)(x-3)=0 because i=sqrt(-1) x^3-2x^2-9x+18=0

OpenStudy (anonymous):

ure a lifesaverrrr thanks soooo much

OpenStudy (anonymous):

You are welcome:)